OpenStudy (anonymous):

What is the weight (in grams) of a liquid that exactly fills a 465.0 milliliter container if the density of the liquid is 0.982 grams over milliliter? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@sweetburger

OpenStudy (danjs):

mass, in grams you still working on this?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

hi

OpenStudy (danjs):

The thing to remember is the definition of density.. The mass of something per a certain unit of volume Density = Mass / volume

OpenStudy (danjs):

A volume of air will be much less dense than the same volume of rocks

OpenStudy (anonymous):

Yes.

OpenStudy (danjs):

So write down given - find -

OpenStudy (danjs):

the givens are - Volume, and Density, Find-- the mass

OpenStudy (anonymous):

Okay, then?

OpenStudy (danjs):

WIthout the paragraph, this is the prob V = 465.0 mL D = 0.982 g/mL m = ?? D = m/V

OpenStudy (anonymous):

So how do I get m?

OpenStudy (danjs):

Solve the density definition for what you want, mass first, then use the values D = mass / Volume mass = density * Volume

OpenStudy (danjs):

Just remember that the density of something is the mass over a volume, D = m/V

OpenStudy (danjs):



OpenStudy (anonymous):

So, which numbers do I multiply?

OpenStudy (danjs):



OpenStudy (anonymous):

Okay, that's what i thought. So, 0.211?

OpenStudy (anonymous):

@DanJS